OpenStudy (anonymous):

In the conversion table: The compound rate is annually = 1(#compounding periods per year), semi-annually = 2, Quarterly = 4, Monthly = 12. Considering all of that, if the compounding rate is 10 years, what is the number of compounding periods?

OpenStudy (experimentx):

what??? compounding rate is 10 years, i believe you will get interest at 10 years

OpenStudy (ash2326):

For annually it's \(1 \times 10=10\) For semi annually \(2 \times 10=20\) now you can find for the rest???

OpenStudy (ash2326):

In a year there will be 2 compounding periods, so in 10 years==> 20 compounding periods

OpenStudy (experimentx):

I think the time is 10 years ... not compounding period. for semi annual compounding period, you will have 20 periods in 10 years, i.e. for each year .. tow compounding period.

OpenStudy (ash2326):

Is this your formula \[FV=R \times \frac{((1+i)^n-1)}{i}\]

OpenStudy (experimentx):

Mistyped ... what is FV by the way??

OpenStudy (ash2326):

But for compound interest the formula is \[FV=R (1+\fr ac{i}{n})^nt\] n= number of times the interest is compounded t= no of years

OpenStudy (ash2326):

\[FV=R(1+\frac{i}{n})^{nt}\]

OpenStudy (ash2326):

Ok got it:)

OpenStudy (experimentx):

i think you are error in your equation FV = R × ((1 + i)n − 1)/i why are you dividing by i ??

OpenStudy (experimentx):

and also (1 + i)n − 1 ---> here subtracting -1 means you are calculating only interest, not the total amount.

OpenStudy (experimentx):

No i don't think so .. see ash's formula

OpenStudy (ash2326):

sorry n=20

OpenStudy (experimentx):

which formula?? dividing by i?? i don't think .. FV must be something else.